the length of a retangle is 4 feet less than 3 times its width. the perimeter is 16 feet. what is its length.
Answer by Rent-A-Whiz(1)

Good Afternoon,
In response to your question:
“the length of a rectangle is 4 feet less than 3 times its width. the perimeter is 16 feet. what is its length.”
Step one: Write the sentence in equation form. Lets call the width W and the length L
If it’s easier translate one part of the sentence at a time starting with the multiplication first. So the first sentence now reads:
“The length of a rectangle is 4 feet less than 3W.”
Now translate the rest:
L=3W-4
For the second sentence you need the formula for the perimeter W+L=perimeter= 16 feet.
Now there are 2 equations with two unknowns:
L=3W-4 (1)
2(W+L)=16 (2)
Substitute for L in equation (2):
W + (3W-4) =16,
Which simplifies to:
W=3
Substitute for this value of W in either equation (1) or (2). If I put it in equation (1) I get:
L=3(3) -4=5
Check results:
Substitute for L=5 and W=4 into equation (2) for the perimeter:
2(3+5)=16
Since the left-hand-side is equal to the right-hand-side this is the correct answer.
Good Luck,
Selma from Rent-A-Whiz. A new online tutoring and homework help resource. For more details visit us at www.rentawhiz.com.